# Classification of Semigraphical Translators

### v1

@@ 5. Nonexistence of The Yeti
$\nabla f(x,y)=v$
for all $x$. If $|y|\leq\pi/4$, then $|f_{y}(x,y)|\leq 1$.
@@
$f(x,y)=c^{2}\log(\cos(y/c))\pm x\sqrt{c^{2}-1},$
where $c=2a/\pi\geq 1$. Thus $f_{x}=\pm\sqrt{c^{2}-1}$, so we can choose $c$ and the sign of $\pm$ so that $f_{x}\equiv v_{1}$. Note that
$f_{y}(x,y)=-c\tan(y/c),$
which ranges from $\infty$- ∞ to +$\infty$+ ∞, so (given $c$) we can choose $y$ so that $f_{y}(x,y)=v_{2}$.
Now $|c\tan(y/c)|$ is a decreasing function of $c\in[1,\infty)$, so, for $|y|\leq\pi/4$,
$|f_{y}(x,y)|$ $=|c\tan(y/c)|$
@@
By definition, a yeti is (up to translation and rotation) a complete translator $M$ such that $Z\subset M$ and such that $M\setminus Z$ is the graph of a function over $\mathbf{R}^{2}\setminus\{y=0\}$. If $u$ is the restriction of that function to the halfplane $\{y>0\}$, then (by completeness) $u$ has boundary value +$\infty$+ ∞ on one component of $\{y=0\}\setminus\{(0,0)\}$ and $\infty$- ∞ on the other component. The following theorem proves that such translators do not exist.
@@
A priori, the subsequential limit might not exist, or it might be a graph over a proper subset of $\mathbf{R}\times(0,\infty)$. To exclude those possibilities, we need suitable gradient bounds on $u$. By [graphs-himw]*Theorem 6.7, to get such bounds, it suffices to show that $|Du|$ is bounded on some connected, unbounded set contained in a halfstrip of the form $[x_{1},\infty)\times I$, where $I$ is compact interval in $(0,\infty)$. Fortunately, we are able to produce the necessary set, namely the graph of a certain smooth function $Y:[x_{1},\infty)\to I$ such that, on that graph, $u_{y}=0$ and $|u_{x}|$ is bounded.
@@
be a function whose graph is a complete translator. (Thus $f=-\infty$ on the boundary.) Then the infimum of $u-f$ on the half-strip
@@
Let $\mathcal{R}_{\theta}:\mathbf{R}^{2}\to\mathbf{R}^{2}$ be counterclockwise rotation through $\theta$ about the point $(-1,0)$. For $\theta\in(0,\pi/2)$, let
@@
Thus $S(\theta)$ is a triangular region with $(-1,0)$ as one vertex. Let $E(\theta)$ be the interior of the edge joining $(-1,0)$ to the vertex in $\{y>0\}$, and let $E(0)$ be its limit as $\theta$ tends 00:
@@
has boundary value +$\infty$+ ∞ at all boundary points except those in the edge $E(\theta)$. By the maximum principle, the minimum of $u-f_{\theta}$ is attained at a point in $E(\theta)$:
@@
There is a $c\in(0,\infty)$ with the following property. For every $x>0$, there is a $y\in(0,c)$ such that $u_{y}(x,y)<0$.
@@
$\liminf_{y\to 0}u_{y}(-1,y)=-\infty.$
Thus we can choose a point $p$ with $x(p)=-1$ such that
[…]
@@
(7) $g_{y}<0$ on $\mathbf{R}\times[0,c)$.
Choose $a$ large enough that $[b,c]\subset(-a,a)$, and let
@@
be a tilted grim reaper function such that $f_{x}<0$. By Claim 1, and by adding a constant to $f$, we can assume that
[…]
Since $[b,c]\subset(-a,a)$, we see that $f_{x}-g_{x}$ is a constant $-k<0$. Thus
@@
Let $h=u-g$. Thus the domain of $h$ is the strip
@@
Since $Dh(p)=0$, $C$ is a tree with at least $4$ ends. From the boundary values of $h$, we see that each end of $C$ tends to $-\widehat{\infty}$, to $\widehat{\infty}$, or to $(0,0)$. Here $\pm\widehat{\infty}$ are the points at $x=\pm\infty$ given by (2).
By (8), no end tends to $-\widehat{\infty}$.
@@
Thus, at least three ends of $C$ tend to $+\widehat{\infty}$. Consequently, for each $\hat{x}>0$, there are at least $3$ values of $y\in[0,c)$ for which $g(\hat{x},y)=0$. Let $0<y_{1}<y_{2}<c$ be two of those values. Then
$u(\hat{x},y_{1})-g(\hat{x},y_{1})=u(\hat{x},y_{2})-g(\hat{x},y_{2}).$
@@
$u(\hat{x},y_{2})-u(\hat{x},y_{1})=g(\hat{x},y_{2})-g(\hat{x},y_{1})<0$
since $g_{y}(x,y)<0$ for $y\in[0,c)$. Thus there is a $y\in(y_{1},y_{2})\subset(0,c)$ such that $u_{y}(x,y)<0$. ∎
@@
Let $c$ be as in Claim 2. Let $Q$ be the set of $(x,y)\in\mathbf{R}\times(0,c)$ such that $u_{y}(x,y)=0$. There is an $x_{1}\in(0,\infty)$ such that $Q\cap\{x\geq x_{1}\}$ is the graph of a smooth function $Y:[x_{1},\infty)\to(\pi/4,c)$. Furthermore,
@@
converges to a complete translator that lies in the halfspace $\{y\geq 0\}$. Let $M$ be the component that contains $p:=(0,\hat{y},0)$.
Note that the $y$-axis is tangent to $M$ at $p$. Thus, $M$ cannot be a vertical plane (since $M$ lies in $\{y\geq 0\}$). Hence, $M$ is a graph. By the classification of graphical translators [graphs-himw], it is the graph of a function $\phi$ over a strip in $\{y\geq 0\}$ of width $\geq\pi$. Also, by that classification, $\phi_{yy}<0$ at all points, and $\phi_{y}=0$ only on the midline of the strip. Since $\phi_{y}(0,\hat{y})=0$, we see that the line $\{y=\hat{y}\}$ is the midline of the strip. Since the strip has width $\geq\pi$,
@@
Thus, we have shown that for every sequence $p_{n}\in Q$ with $x(p_{n})\to\infty$,
@@
    every subsequential limit of $u_{yy}(p_{n})$ is $<0$.
@@
Let $x\geq x_{1}$. We claim there is exactly one $y\in(0,c)$ for which $u_{y}(x,y)=0$, i.e., for which $(x,y)\in Q$.
Now $c$ was chosen according to Claim 2. Thus there is a $\tilde{y}\in(0,c)$ for which $u_{y}(x,\tilde{y})<0$. Since $u(x,0)=-\infty$, there is at least one value of $y\in(0,\tilde{y})$ for which $u_{y}(x,y)=0$.
@@
would attain its minimum at some $\hat{y}\in(y_{1},y_{2})$. Thus $u_{y}(x,\hat{y})=0$ and $u_{yy}(x,\hat{y})\geq 0$, contrary to (10).

### v2

@@ 5. Nonexistence of The Yeti
$Df(x,y)=v$
for all $x$. If $|y|\leq\pi/4$, then $|\frac{\partial f}{\partial y}(x,y)|\leq 1$.
@@
$f(x,y)=\kappa^{2}\log(\cos(y/\kappa))\pm x\sqrt{\kappa^{2}-1},$
where $\kappa=2a/\pi\geq 1$. Thus $\frac{\partial f}{\partial x}=\pm\sqrt{\kappa^{2}-1}$, so we can choose $\kappa$ and the sign of $\pm$ so that $\frac{\partial f}{\partial x}\equiv v_{1}$. Note that
$\frac{\partial f}{\partial y}(x,y)=-\kappa\tan(y/\kappa),$
which ranges from -$\infty$ to $\infty$, so (given $\kappa$) we can choose $y$ so that $\frac{\partial f}{\partial y}(x,y)=v_{2}$.
Now, given $|y|\leq\pi/4$, we know that $\kappa\mapsto|\kappa\tan(y/\kappa)|$ is a decreasing function of $\kappa\in[1,\infty)$, so
$\left|\frac{\partial f}{\partial y}(x,y)\right|$ $=|\kappa\tan(y/\kappa)|$
@@
By definition, a yeti is (up to translation and rotation) a complete translator $M$ such that $Z\subset M$ and such that $M\setminus Z$ is the graph of a function over $\mathbf{R}^{2}\setminus\{y=0\}$. If $u$ is the restriction of that function to the halfplane $\{y>0\}$, then (by completeness) $u$ has boundary value $\infty$ on one component of $\{y=0\}\setminus\{(0,0)\}$ and -$\infty$ on the other component. The following theorem proves that such translators do not exist.
@@
Remark 9.
The hypothesis that $\operatorname{graph}(u)\cup Z$ is a smooth manifold-with-boundary is not necessary. See Proposition 11.
@@
A priori, the subsequential limit might not exist, or it might be a graph over a proper subset of $\mathbf{R}\times(0,\infty)$. To exclude those possibilities, we need suitable gradient bounds on $u$. By [graphs-himw]*Theorem 6.7, to get such bounds, it suffices to show that $|Du|$ is bounded on some connected, unbounded set contained in a halfstrip of the form $x_{1},\infty)\times I$, where $I$ is compact interval in $(0,\infty)$. Fortunately, we are able to produce the necessary set, namely the graph of a certain smooth function $Y:[x_{1},\infty)\to I$ such that, on that graph, $\frac{\partial u}{\partial y}=0$ and $|\frac{\partial u}{\partial x}|$ is bounded.
@@
be a function whose graph is a complete translator. (Thus $f=$-$\infty$ on the boundary.) Then the infimum of ufu-f on the half-strip
@@
Let $\mathcal{R}_{\theta}:\mathbf{R}^{2}\to\mathbf{R}^{2}$ be counterclockwise rotation through $\theta$ about the point (1,0)(-1,0). For $\theta\in(0,\pi/2)$, let
@@
Thus $S(\theta)$ is a triangular region with (1,0)(-1,0) as one vertex. Let $E(\theta)$ be the interior of the edge joining (1,0)(-1,0) to the vertex in $\{y>0\}$, and let $E(0)$ be its limit as $\theta$ tends 0:
@@
has boundary value $\infty$ at all boundary points except those in the edge $E(\theta)$. By the maximum principle, the minimum of $u-f_{\theta}$ is attained at a point in $E(\theta)$:
@@
There is a $c\in(0,\infty)$ with the following property. For every $x>0$, there is a $y\in(0,c)$ such that $\frac{\partial u}{\partial y}(x,y)<0$.
@@
$\liminf_{y\to 0}\frac{\partial u}{\partial y}(-1,y)=-\infty.$
Thus we can choose a point $p$ with x(p)=1x(p)=-1 such that
(7) $\begin{gathered}\text{$0<y(p)<\pi/4$, and}\\ \frac{\partial u}{\partial y}(p)<-1.\end{gathered}$
@@
(8) $\frac{\partial g}{\partial y}<0$ on $\mathbf{R}\times[0,c)$.
Choose $a$ large enough that $b,c]\subset(-a,a)$, and let
@@
be a tilted grim reaper function such that $\frac{\partial f}{\partial x}<0$. By Claim 1, and by adding a constant to $f$, we can assume that
$\text{$u-f\geq 0$ on $(-\infty,-1]\times(0,c)$}.$
Since $b,c]\subset(-a,a)$, we see that $\frac{\partial f}{\partial x}-\frac{\partial g}{\partial x}$ is a constant k<0-k<0. Thus
@@
Let h=ugh=u-g. Thus the domain of $h$ is the strip
@@
Since $Dh(p)=0$, $C$ is a tree with at least $4$ ends. From the boundary values of $h$, we see that each end of $C$ tends to $\widehat{\infty}$, to $\widehat{\infty}$, or to $(0,0)$. Here $\pm\widehat{\infty}$ are the points at $x=\pm\infty$ given by (3).
By (9), no end tends to $\widehat{\infty}$.
@@
Thus, at least three ends of $C$ tend to $\widehat{\infty}$. Consequently, for each $x>0$, there are at least $3$ values of $y\in[0,c)$ for which $h(x,y)=0$. Let $0<y_{1}<y_{2}<c$ be two of those values. Then
$u(x,y_{1})-g(x,y_{1})=u(x,y_{2})-g(x,y_{2})=0.$
@@
$u(x,y_{2})=g(x,y_{2})<g(x,y_{1})=u(x,y_{1}),$
since $\frac{\partial g}{\partial y}(x,y)<0$ for $y\in[0,c)$. Thus there is a $y\in(y_{1},y_{2})\subset(0,c)$ such that $\frac{\partial u}{\partial y}(x,y)<0$. ∎
@@
Let $c$ be as in Claim 2. Let $Q$ be the set of $(x,y)\in\mathbf{R}\times(0,c)$ such that $\frac{\partial u}{\partial y}(x,y)=0$. There is an $x_{1}\in(0,\infty)$ such that $Q\cap\{x\geq x_{1}\}$ is the graph of a smooth function $Y:[x_{1},\infty)\to(\pi/4,c)$. Furthermore,
@@
converges to a complete translator that lies in the halfspace $\{y\geq 0\}$. Let $\widehat{M}$ be the connected component of the limit that contains $p:=(0,\hat{y},0)$.
Note that the $y$-axis is tangent to $\widehat{M}$ at $p$. Thus, $\widehat{M}$ cannot be a vertical plane (since $\widehat{M}$ lies in $\{y\geq 0\}$). Hence, $\widehat{M}$ is a graph. By the classification of graphical translators [graphs-himw], it is the graph of a function $\phi$ over a strip in $\{y\geq 0\}$ of width $\geq\pi$. Also, by that classification, $\frac{\partial^{2}\phi}{\partial y^{2}}<0$ at all points, and $\frac{\partial\phi}{\partial y}=0$ only on the midline of the strip. Since $\frac{\partial\phi}{\partial y}(0,\hat{y})=0$, we see that the line $\{y=\hat{y}\}$ is the midline of the strip. Since the strip has width $\geq\pi$,
@@
Thus, we have shown that every sequence $p_{n}\in Q$ with $x(p_{n})\to\infty$ has a subsequence $p_{n(i)}$ such that
@@
    $\frac{\partial^{2}u}{\partial y^{2}}(p_{n(i)})$ converges to a limit $\frac{\partial^{2}\phi}{\partial y^{2}}(0,\hat{y})<0$.
@@
Let $x\geq x_{1}$. We claim there is exactly one $y\in(0,c)$ for which $\frac{\partial u}{\partial y}(x,y)=0$, i.e., for which $(x,y)\in Q$.
Now $c$ was chosen according to Claim 2. Thus there is a $\tilde{y}\in(0,c)$ for which $\frac{\partial u}{\partial y}(x,\tilde{y})<0$. Since $u(x,0)=$-$\infty$, there is at least one value $y\in(0,\tilde{y}) fo$r which $\frac{\partial u}{\partial y}(x,y)=0$.
@@
would attain its minimum at some $\hat{y}\in(y_{1},y_{2})$. Thus $\frac{\partial u}{\partial y}(x,\hat{y})=0$ and $\frac{\partial^{2}u}{\partial y^{2}}(x,\hat{y})\geq 0$, contrary to (11).
@@
The following proposition was not used in this paper, but it justifies the assertions in Remarks 4, 6, and 9.
Proposition 11.
Suppose that
$u:\{(x,y)\in B^{2}(0,R):y>0\}\to\mathbf{R}$
is a smooth solution to the translator equation such that
$u(x,0)=\begin{cases}-\infty&\text{if $x<0$},\\ \infty&\text{if $x>0$}.\end{cases}$
Then $\operatorname{graph}(u)\cup Z$ is a smooth-manifold-with-boundary.
Proof.
Let $C(R)$ be the cylinder $B^{2}(0,R)\times\mathbf{R}$. Let $M$ be the graph of $u$. Then $M$ minimizes the translator area in the following sense: if $K$ is a compact portion of $M$ and if $K^{\prime}$ is a compact surface in $C(R)\cap\{y>0\}$ with $\partial K^{\prime}=\partial K$, then
$\operatorname{area}(K)\leq\operatorname{area}(K^{\prime}).$
Here and in what follows, we use $\operatorname{area}(\Sigma)$ to refer to the translator area of a surface $\Sigma$, computed with respect to Ilmanen’s translator metric $g_{ij}(x,y,z)=e^{-z}\delta_{ij}$.
Actually, the assumption that $K^{\prime}$ lies in $C(R)\cap\{y>0\}$ is not needed. For suppose that $K^{\prime}$ is any compact surface with $\partial K^{\prime}=\partial K$. Let $Q$ be a compact convex subset of $B^{2}(0,R)\cap\{y>0\}$ such that $Q\times\mathbf{R}$ contains $K$ and let $K^{\prime\prime}$ be the projection of $K^{\prime}$ to $Q\times\mathbf{R}$. Then
$\operatorname{area}(K)\leq\operatorname{area}(K^{\prime\prime})\leq\operatorname{area}(K^{\prime}).$
In particular, if $K$ is a convex, compact subset of $C(R)\cap\{y>0\}$, then $M\cap K$ and $V\cap\partial K$ have the same boundary, where
$V=\{(x,y,z)\in C(R):y>0,\,z<u(x,y)\}$
is the subgraph of $u$. Thus
$\operatorname{area}(M\cap K)$ $\leq\operatorname{area}(V\cap\partial K)$
$\leq\operatorname{area}(\partial K)$
Thus, if $K$ is any compact, convex subset of $C(R)$, then
$\operatorname{area}(M\cap K\cap\{y\geq\epsilon\})$ $\leq\operatorname{area}(\partial K\cap\{y\geq\epsilon\})$
$\leq\operatorname{area}(\partial K).$
Thus, letting $\epsilon\to 0$,
$\operatorname{area}(M\cap K)\leq\operatorname{area}(\partial K).$
The fact that $M\cap K$ has finite area together with the fact that $M\cap\{y=0\}$ has area 0 implies that the surfaces
($\star$) $M\cap K\cap\{y>\epsilon\}$
converge (as currents, with the mass norm) to $M\cap K$ as $\epsilon\to 0$. Thus, $M\cap K$ is area-minimizing, since it is the limit of the area-minimizing surfaces ($\star$5).
Let $S$ be the strip
$S=\{(x,0,z):x\in(0,R),\,z\in\mathbf{R}\}.$
Let [[M]][\![M]\!], [[S]][\![S]\!], and [[V]][\![V]\!] be the locally integral currents obtained by orienting $M$ by the upward unit normal, and $S$ by the unit normal $\mathbf{e}_{2}$, and by giving $V$ the standard orientation.
Then
[…]
so
[…]
$=[\![Z]\!].$
We have therefore proved that $M$ is translator-area minimizing and has boundary (in $C(R)$) the line $Z$ with multiplicity $1$. Hence by [hardt-simon], $M\cup Z$ is a smooth manifold-with-boundary. ∎